Cindy distributed 150 sweets among five bottles. Bottle A and B together contain 58 sweets. Bottle B and Bottle C together contain 45 sweets,Bottle C and Bottle D together contain 69 sweets and Bottle D and Bottle E together contain 59 sweets.How many sweets are there in each bottle? With exact solution.
Answer by ankor@dixie-net.com(22740)

Write an equation for each statement
:
Cindy distributed 150 sweets among five bottles.
A + B + C + D + E = 150
:
Bottle A and B together contain 58 sweets.
A + B = 58
:
Bottle B and Bottle C together contain 45 sweets
B + C = 45,
:
Bottle C and Bottle D together contain 69 sweets
C + D = 69
:
and Bottle D and Bottle E together contain 59 sweets.
D + E = 59
:
Back to the first equation
A + B + C + D + E = 150
Replace A + B with 58; replace D + E with 59
58 + C + 59 = 150
117 + C = 150
C = 150 - 117
C = 33 candies
Find D, replace C with 33
33 + D = 69
D = 69 - 33
D = 36
In the same manner, You should be able to find A, B and E using the given equations
Total up all 5 values, to confirm they add up to 150
:
:
:
How many sweets are there in each bottle? With exact solution.